A 22.7 mL sample of 0.318 M dimethylamine, (CH3)2NH, is titrated with 0.257 M nitric acid. At the titration midpoint, the pH is ____.
Answer #1

(CH3)2NH + HNO3 -------------> (CH3)2NH2+ + NO3-

millimoles of base = 22.7 x 0.318 = 7.22

7.22 millimoles HNO3 must be added to reach equivalence point.

7.22 = V x 0.257

V = 28.1 mL

total volume = 22.7 + 28.1 = 50.8 mL

[salt] = 7.22 / 50.8 = 0.142 M

for salt of weak base

pOH = 1/2 [pKw + pKb + logC]

pKb = 3.27 standard value

pOH = 1/2 [14 + 3.27 + log 0.142]

pOH = 8.21

pH = 14 - 8.21

pH = 5.79